### v2

# Super-Resolution with Structured Motion

Abstract
We consider the limits of super-resolution using imaging constraints. Due to various theoretical and practical limitations, reconstruction-based methods have been largely restricted to small increases in resolution. In addition, motion-blur is usually seen as a nuisance that impedes super-resolution. We show that by using high-precision motion information, sparse image priors, and convex optimization, it is possible to increase resolution by large factors. A key operation in super-resolution is deconvolution with a box. In general, convolution with a box is not invertible. However, we obtain perfect reconstructions of sparse signals using convex optimization. We also show that motion blur can be helpful for super-resolution. We demonstrate that using pseudo-random motion it is possible to reconstruct a high-resolution target using a single low-resolution image. We present numerical experiments with simulated data and results with real data captured by a camera mounted on a computer-controlled stage.
Index Terms:
Computational Photography, Super-resolution, Compressed Sensing, Deconvolution
1 Introduction
We consider the super-resolution problem where the goal is to reconstruct a high-resolution image from one or several low-resolution images. We focus primarily on the use of imaging constraints and low-level image priors, and on the use of motion to capture one or more pictures.
One of our motivations is to implement a compressed sensing system for super-resolution. Using a high-resolution CCD/CMOS sensor, it may be possible to reconstruct an extremely high-resolution image using an appropriate optical and computational system.
Classical methods for super-resolution relate a collection of low-resolution images to a high-resolution image using a model of the imaging process. Such methods have been largely restricted to relatively small increases in resolution due to theoretical and practical limitations [1, 2]. In addition, motion blur is usually seen as a nuisance that impedes super-resolution [3].
Despite the theoretical limits previously described in the literature, we show that it is possible to increase resolution by large factors using sparse image priors. We note a connection between super-resolution and deconvolution with a box filter, which helps to understand the ambiguities in the super-resolution problem. In particular, a collection of low-resolution images determines most of the Fourier components of the super-resolution image and, in this setting, sparse images can be recovered using convex optimization.
We also show that by using motion blur, it is possible to reconstruct a high-resolution target from a single low-resolution image. By moving the camera or sensor during an exposure, we capture a superposition of multiple low-resolution images. Figure 1 illustrates the idea. In this example where we capture an image while the camera/sensor undergoes a large vibration. We record a single low-resolution image and reconstruct a high-resolution target by solving a convex optimization problem.
Refer to caption Refer to caption Refer to caption Refer to caption
Ground truth Sensor trajectory Recorded image Reconstruction
$128\times 128$ $128\times 128$ $64\times 64$ $128\times 128$
Figure 1: Using motion blur for super-resolution.
Super-resolution imaging systems have many potential applications. In some settings, there are practical and physical limitations that lead to low-resolution sensors. Furthermore, there are many applications where a camera might be constantly moving, such as in aerial imaging. High-resolution imaging has many applications in science and in archival/preservation work (including for art, documents, and biological samples). For example, [4] describes a large format tile-scan camera to image static scenes for cultural-heritage preservation.
For the case of visible light photography with high-resolution sensors, our methods could be used to produce gigapixel images using a small camera. Pixel shift cameras are commercially available and already use sensor motion to overcome the resolution limit induced by color filters. The same technology could be used to implement the imaging methods described here.
We implemented a system to demonstrate some of our ideas using a commonly available camera and computer-controlled motion stage. The experiments in Section 6 demonstrate results using both static images and images taken with the camera moving at a constant velocity.
Throughout the paper, we focus on cases where the effective resolution is limited by the sensor and not by the optical path (diffraction). Even with high-resolution sensors, super-resolution can be useful for obtaining a wider field of view with an appropriate lens. The interlacing operation described in Section 6.2 is reminiscent of a panorama, but interlaces a set of pictures taken with small motion instead of tiling pictures taken with large motion, which leads to a more compact design for a high-resolution camera.
We assume throughout the paper that the sensor motion is known, either because it is carefully planned and executed by a high-accuracy actuator or because we have a high-accuracy readout of the sensor position while it is being moved through some other means (such as vibration). In practice, the motion information should be accurate to the target pixel resolution. In Section 6.3, we use a landmark (sharp edge) to estimate a high-resolution camera position directly from a low-resolution image.
2 Related Work
The super-resolution problem has a long history (see, e.g.[5, 6, 1, 2]). From a theoretical point of view, the results in [1] and [2] are largely negative, showing that ambiguity in the super-resolution problem grows quickly with the desired increase in resolution. In contrast, here we note that with a particular sampling scheme, this ambiguity is concentrated on relatively few Fourier coefficients. Previous work in compressed sensing [7, 8] has shown that sparse image priors can be used to resolve such ambiguities.
[3] proposed the use of a ”Jitter camera” to improve the resolution of video by quickly moving the sensor in the imaging plane of a camera without introducing motion blur, which was believed to impede super-resolution. Here we argue that motion blur can in fact aid in super-resolution.
Intuitively, the combination of motion blur and discrete sampling can be used to define a compressed sensing system. The approach is related to the use of random filters for compressed sensing [9, 10].
A variety of unique imaging system designs have been developed through the use of compressed sensing methods. Some physical implementations include [11, 12, 13].
Recently, some unusual imaging schemes have been proposed, not to achieve super-resolution, but to remove motion-related artifacts. For example, parabolic lens motion [14], induced by constant acceleration, generates a PSF that is invariant to object velocity. Additionally, fluttering a camera shutter [15, 16] has been shown to help in removing motion blur. This leads to a surprising idea that it may be beneficial to extend an exposure with a coded shutter to compensate for motion blur.
Many super-resolution methods use training data or other sources of information to improve image resolution. The idea of exemplar based super-resolution goes back to the work in [1] and was implemented at a large scale in [17]. Generic image priors based on self-similarity have also been used for super-resolution [18]. More recent methods have used deep neural networks for super-resolution from a single image[19].
In this paper, we focus on the use of total variation for reconstruction. In practice, alternative image priors, such as ones based on self-similarity or deep networks, could also be implemented using Plug-and-Play methods [20]. Total variation was also used as prior for super-resolution in [6].
Neural models have been increasingly used for image reconstruction, including super-resolution [21]. The use of a neural radiance field (NeRF) allows for the estimation of a continuous image instead of selecting a particular target resolution. In this case, one can implement the imaging model described here by Monte Carlo integration via sampling a finite number of locations in the imaging plane.
We consider only grayscale images to simplify the presentation and experimental design. For the case of color images captured with a color filter array, one could reconstruct a super-resolution image for each color channel separately, using measurements defined by pixels of that color in the filter array. This is related to the recent approach in [22] which performs super-resolution without explicit demosaicing. By using controlled motion, one can also collect measurements for every pixel using each color in the filter array.
3 Imaging Model
Our goal is to reconstruct a high-resolution image from one or several low-resolution images. In the mathematical model we consider here, the low-resolution images are captured with a sensor that can translate in the imaging plane of a camera. In one setting, we capture a series of images in a grid of high-resolution (sub-pixel) locations. In another setting, we capture one or more images while the sensor is moving; for example, we can record a single or sequence of pictures while vibrating the sensor or moving along a linear trajectory (e.g, as may be the case in satellite imaging).
There are many equivalent practical situations. In the experiments in Section 6, we use a standard camera on a moving stage, which is equivalent to moving the sensor when viewing a planar object parallel to the camera. In a microscope, instead of moving the sensor, we can precisely move the sample using a piezo nanopositioning stage.
Let $g:\mathbb{R}^{2}\rightarrow\mathbb{R}$ denote a continuous brightness function in the imaging plane of a camera. We do not model the optics of the camera and focus only on measuring the image projected onto the imaging plane.
An $n\times m$ sensor integrates $g$ over square pixels tiling a rectangular area. We use $\Delta$ to denote the side length of a pixel in the sensor. Each pixel integrates the light that falls in a particular square region over a unit of time.
Ignoring measurement noise, if we place the sensor in the imaging plane with bottom-left corner at $(x_{0},y_{0})$ we capture a discrete image $I$ with
[…]
We will also consider the case where the sensor moves while the image is captured. If the sensor follows a trajectory $p(t)=(x_{0}(t),y_{0}(t))$, we capture an image $I$ with
[…]
Note that translating the sensor by $z$ is equivalent to translating $g$ by $-z$. Let $q(x,y)$ denote the occupancy map defined by the trajectory $-p(t)$. That is, $q(x,y)$ is the fraction of time the trajectory spends at position $(x,y)$. The image $I$ captured by the moving sensor is equivalent to the image of $g\otimes q$ captured by a static sensor.
We consider methods that increase the resolution of the camera sensor to generate a super-resolution image $J$ that would have been recorded by a static sensor with pixels of size $\Delta/f$ for an integer factor $f$ defining the resolution increase. This virtual sensor covers an area that is at least as big as the camera sensor and the image $J$ has at least $fn\times fm$ pixels (the area may be larger due to the sensor motion). In our experiments, we have used super-resolution factors as high as $f=8$.
4 Grid of Images
Let $J$ be an image generated by a (virtual) sensor with pixel size $\Delta/f$ at the origin. Let $I$ be an image captured by a (physical) sensor with pixel size $\Delta$ at $(x_{0},y_{0})$ where
$x_{0}=k(\Delta/f),$
$y_{0}=l(\Delta/f),$
with $k$ and $l$ integers. In this case, the low-resolution pixels in $I$ are aligned with the high-resolution pixels in $J$. Therefore, each value in $I$ is the sum of $f^{2}$ values in $J$,
$I[i,j]=\sum_{i^{\prime}=0}^{f-1}\sum_{j^{\prime}=0}^{f-1}J[k+fi+i^{\prime},l+fj+j^{\prime}].$
We see that $I$ is a decimation (subsampling) of $J\otimes B$ where $B$ is a two-dimensional box filter of width $f$.
Now suppose we capture $f^{2}$ images $\{I_{k,l}\}$ by translating the sensor over an $f\times f$ grid of high-resolution locations,
[…]
Capturing such images involves moving the sensor to sub-pixel locations defined by a step size $\Delta/f$. The total vertical and horizontal displacement required to capture all images is less than $\Delta$.
4.1 Interlacing
The images $\{I_{k,l}\}$ can be rearranged into a single $fn\times fm$ image $H$ such that $H=J\otimes B$, where $B$ is a box of width $f$. Since each image $I_{k,l}$ is a decimation of $J\otimes B$, the rearrangement interlaces the captured images:
$H[k+if,l+jf]=I_{k,l}[i,j].$
Figure 2 illustrates the process where we capture multiple low-resolution images at sub-pixel shifts and interlace the result to obtain a single image $H$. The interlaced image $H$ has $fn\times fm$ pixels and defines the imaging constraints on $J$ through the relation $H=J\otimes B$.
Figure 2: Interlacing $f\times f$ low-resolution images taken in a grid of sub-pixel locations, we obtain the convolution of the high-resolution target with a two-dimensional box filter of width $f$.
4.2 Deconvolution with a Box
Recovering the super-resolution image $J$ is equivalent to a deconvolution of $H$ with a two-dimensional box filter of width $f$. That is, we need to de-blur $H$ to recover $J$.
Note that a box filter is not a good smoothing or low-pass filter. Convolution with a box yields a signal that is visually blurry but still has quite a lot of high-frequency information. In Figure 3(a) we compare the Fourier transform of a one-dimensional box and a Gaussian of similar width. The Fourier transform of a box is a sinc, and the Fourier transform of a Gaussian is a Gaussian. Although the sinc has isolated zeros, the magnitude of the Fourier transform of the box decays much slower than the magnitude of the Fourier transform of a Gaussian.
For a two-dimensional discrete (and periodic) signal, we illustrate the magnitude of the DFT of a box of width $4$ in Figure 3(b). Note that there are relatively few coefficients in the DFT of $B$ with magnitude close to 0. We can therefore directly recover most of the coefficients in the Fourier transform of $J$ from the Fourier transform of $H$, except for isolated frequencies where the Fourier transform of $B$ is zero or close to zero.
Figure 3: (a) One-dimensional box and Gaussian (top) and their Fourier transforms (bottom). The red plots correspond to a Gaussian and the blue plots correspond to a box. (b) Bode plot (magnitude of Fourier transform in dB) of a two-dimensional discrete box filter of width 4.
This is exactly the setting in which sparse image priors and convex optimization are known to yield surprisingly good (“perfect”) reconstructions [7, 8].
For one-dimensional sparse signals, deconvolution with a box using $\ell_{1}$ regularization leads to perfect reconstructions [23]. This is true despite the ambiguity in the imaging constraints due to the unknown Fourier coefficients. The key here is the choice of prior which selects a particular solution among all solutions that agree (or nearly agrees in the case of noisy data) with the measurements.
Figure 4 illustrates a numerical experiment with a simulated camera. The deconvolution of $H$ using a total variation prior yields an almost perfect reconstruction including very high-resolution details.
(a) Input ($J$)
(b) 1 of 64 low-res images ($I_{k,l}$)
(c) Interlaced images ($H$)
(d) Wiener deconvolution
(e) TV deconvolution
Figure 4: Super-resolution with $f=8$. By interlacing 64 low-resolution images, we obtain the measurements $H=J\otimes B$. Deconvolving $H$ with a TV prior leads to an almost perfect reconstruction.
4.3 The Limits of Super-Resolution
Previous work [1] considered the setting discussed here and showed that the ambiguity in the super-resolution problem grows quickly with the super-resolution factor $f$. This was characterized in terms of the volume of the set of solutions that are compatible with a set of images and the condition number of a resulting linear system.
We note, however, that the negative results in [1] do not account for the use of a prior to select a particular solution among the ones that are compatible (or nearly compatible) with the measured data. For the experiments with imaging constraints in [1] the ambiguity was resolved with a quadratic smoothness prior, while we use total variation.
The difference between using a quadratic smoothness prior and total variation may seem subtle at first degree, but one of the most striking results in compressed sensing is that priors that induce sparsity, like total variation, can lead to perfect reconstructions with under-determined linear systems [7]. The fact that natural images have the appropriate sparsity structure is a surprising phenomenon of its own. Total variation is also known to preserve/recover sharp edges in image restoration (see, e.g. [24]).
Figure 5 compares a result that minimizes a quadratic smoothness prior (as in [1]) to a result that minimizes the total variation. In this case there is no measurement error and both solutions satisfy the imaging constraints defined by $H=J\otimes B$. This illustrates how the solution is ambiguous (as expected) and how the choice of image prior can have a significant impact on the quality of the results. Moreover, the example illustrates that a low-level prior is sufficient to resolve high-resolution details.
(a) Interlaced images ($H$)
(b) Quadratic prior
(c) TV prior
Figure 5: Deconvolution of $H$ with a quadratic smoothness prior leads to an overly smooth image and ringing artifacts while using a TV prior yields an almost perfect reconstruction.
4.4 Wiener Deconvolution
Using a Gaussian prior for the Fourier coefficients of the high-resolution image leads to a very fast Wiener deconvolution algorithm via the FFT and an inverse transform.
The Wiener filter computes a MAP estimate of the super-resolution image $J$. The algorithm is defined by a single parameter $\gamma$ used to shrink the estimated Fourier coefficients.
$\bar{B}$ $=\operatorname*{FFT}(B)$
$\bar{H}$ $=\operatorname*{FFT}(H)$
$\bar{J}$ $=\frac{\bar{H}}{\bar{B}+\gamma}$
$J$ $=\operatorname*{IFFT}(\bar{J})$
4.5 Sparse Reconstruction with TV prior
As discussed above, previous work has shown that sparse signals can often be perfectly or nearly perfectly recovered from incomplete frequency information [7, 8].
In practice, we regularize the reconstruction of the super-resolution image $J$ using the total variation (TV) measure. This leads to a convex optimization problem,
$\min_{J}||H-(J\otimes B)||_{2}^{2}+\lambda TV(J)$
where $TV(J)$ is the total-variation of $J$,
$TV(J)=\sum_{i,j}|J[i,j]-J[i+1,j]|+|J[i,j]-J[i,j+1]|.$
and $\lambda$ is a regularization parameter.
5 Super-resolution by Deblurring
In the previous section, we considered the case of multiple static images taken from different sensor locations. Here we consider the case where one or more images are captured while the sensor is moving.
First, we note that a small amount of motion blur can be helpful for super-resolution. Consider the problem of localizing a point source in one dimension; without motion, we can only localize the point source to the pixel size. Now suppose we move the sensor/camera with constant velocity by exactly one pixel while taking a picture. Figure 6 illustrates this situation. The motion leads to a convolution of the brightness function with a box. The point source is spread over two pixels and we observe two non-zero values $b=I[k]$ and $c=I[k+1]$. Together, these measurements determine both the total light intensity and the sub-pixel point source location,
$a=b+c,$
$t_{0}=(k+0.5)\Delta+c/(b+c)).$
Optical blurring, such as due to diffraction, can lead to a similar effect (see, e.g.[25]) but the use of motion may have advantages since, as discussed in the previous section, convolution with a box preserves more information when compared to a Gaussian filter of similar width.
Figure 6: We can localize a point source to sub-pixel location if the camera undergoes constant velocity motion of one pixel during capture. The motion spreads the source to two pixels, and the relative values determine the source location. On the left, we have the continuous brightness function and the image captured with a static sensor. On the right, the sensor moves with constant velocity by one pixel during capture. The effect of the sensor motion is equivalent to a convolution of the brightness function with a box.
We also note that for “sparse” scenes, we can use large motions to (in essence) take multiple low-resolution images in different areas of the sensor. Figure 7 illustrates a simulation where the sensor undergoes a large magnitude vibration (top row) while capturing an image. This leads to the somewhat counterintuitive notion that shaking the camera during an exposure can lead to sharper pictures.
Being able to take pictures while the sensor/camera is moving allows for faster acquisitions, which is often important. In this case, there is no need to wait for the sensor to move and stabilize in one position before taking a picture; instead, we can follow a continuous trajectory and take a sequence of pictures along the way.
As in the previous section, let $J$ be a super-resolution image generated by a virtual sensor at the origin. Now let $I$ be an image captured by a moving sensor. Let $p(t)=(k(t),l(t))$ for $0\leq t\leq 1$ denote a sensor trajectory on the grid of high-resolution pixel locations.
Let $B$ be a two-dimensional box of width $f$ and $Q$ be the occupancy map of the trajectory $-p(t)$. When the sensor undergoes continuous motion we use linear interpolation to define a discrete high-resolution occupancy map $Q$. The relationship between $J$ and $I$ is captured by a decimation by factor $f$ of the sequential convolution of $J$ with $Q$ and $B$. The convolution of $J$ with $Q$ accounts for the sensor motion, while the convolution with $B$ and decimation accounts for the relative size of the physical and virtual pixels.
Let $(J)\downarrow_{f}$ denote the decimation of $J$ by a factor $f$. The imaging constraints are summarized by the linear system,
$I=(J\otimes Q\otimes B)\downarrow_{f}.$
Due to the decimation by $f$, this system of equations has many fewer constraints than variables. The convolution with $B$ also introduces ambiguities. Nonetheless, these constraints can lead to near perfect reconstructions with an appropriate choice of image prior as we demonstrate in simulations below.
In some applications we may have a sequence of low-resolution images $I_{1},\ldots,I_{s}$ taken over one or multiple trajectories. Let $Q_{j}$ denote the occupancy map of the trajectory followed during the acquisition of $I_{j}$.
We can reconstruct a high-resolution image using the data from multiple motion blurred images simultaneously. To estimate a sparse image using $\ell_{1}$ regularization we solve the convex optimization problem,
$\min_{J}\sum_{k=1}^{s}||I_{k}-(J\otimes Q_{k}\otimes B)\downarrow_{f}||_{2}^{2}+\lambda||J||_{1}.$
Similarly we can regularize the reconstruction using TV,
$\min_{J}\sum_{k=1}^{s}||I_{k}-(J\otimes Q_{k}\otimes B)\downarrow_{f}||_{2}^{2}+\lambda TV(J).$
Here again $\lambda$ is a regularization parameter.
(a) $J$ ($128\times 128$)
(b) static $I$ ($32\times 32$)
(c) trajectory
(d) moving $I$ ($32\times 32$)
(e) Reconstructed $J$
Figure 7: (a) High-resolution targets with 10 random characters. (b) Images recorded by a static low-resolution sensor. (c) Sensor trajectory (vibration on the top and random translations on the bottom). (d) Low-resolution images recorded by a moving sensor. (e) Reconstruction of the high-resolution target from the image recorded by the moving sensor.
5.1 Simulations
To demonstrate the use of motion blur for super-resolution, we simulated images obtained using different sensor trajectories. For the experiments in this section, we treat the high-resolution target as periodic. The sensor motion therefore leads to a cyclic convolution between the target and the sensor occupancy map. We used the $\ell_{1}$ norm of the target image $J$ to regularize the reconstructions.
Figure 7(a) shows two examples of the high-resolution targets used in the experiments. Each target is a $128\times 128$ image with several random characters placed in random locations. Figure 7(b) shows the results of taking a low-resolution picture of each target using a $32\times 32$ sensor. In this case, there is no sensor motion. We can see that with such a low-resolution picture, the characters become completely unreadable.
Figure 7(d) shows the results of imaging the two targets from Figure 7(a) with a moving sensor. In each case, the sensor undergoes a large motion. Figure 7(e) shows the reconstructions of the targets from the low-resolution image captured while the sensor is moving. In these examples, we obtain a nearly perfect high-resolution image from a single low-resolution image by leveraging motion blur.
We consider the following trajectories for the sensor:
  1. 1.
    Vibration (vibration). The $x$ and $y$ coordinates of the moving sensor are defined by two sinusoidals with large magnitude. (Figure 7 top row)
  2. 2.
    1000 random displacements (shifts). In this case, the trajectory is a sequence of 1000 random locations. Each location is selected independently from a uniform distribution in the high-resolution grid. (Figure 7 bottom row)
  3. 3.
    Random walk with 1000 steps (walk). In this case, the sensor undergoes a random walk in the grid of high-resolution locations. (Not shown)
Figure 8 shows the mean RMS error obtained with each trajectory for targets of increasing complexity. We see that for sparse targets, where the number of characters in the image is small, it is possible to obtain perfect 128x128 images using a single 32x32 blurry picture.
Figure 8: Mean reconstruction error for targets with different numbers of characters. Each target is 128x128 and we capture a single 32x32 image undergoing motion.
6 Experimental Results
We implemented the reconstruction algorithms in python. We used numpy for the Wiener deconvolution method and both cvxpy and pylops for computing reconstructions using sparse image priors. We found that using the OSQP solver within cvxpy yields good results with small images, as seen in the simulations throughout the paper. For the experiments with real data, we used the split-bregman algorithm implemented in pylops due to the computational demands of processing larger images.
All of the experiments were done with a desktop computer with a 13th Gen Intel(R) Core(TM) i5-13400F CPU with 16 cores and 32GB of RAM.
When $f=8$, reconstructing a 1200x1520 image from 64 separate 150x190 low-resolution images with the Wiener deconvolution algorithm took about 0.4 seconds. Computing a TV reconstruction using pylops with the same data took between 5 and 10 minutes depending on various tolerances and algorithm parameters.
All of the running times scale linearly with the image sizes. A direct reconstruction of a very large (gigapixel) image using a sparse image prior would be impractical with our current implementation. However, we could break the image into (overlapping) tiles and reconstruct each tile independently in parallel.
6.1 Setup
The data used in the following reconstructions consist of sets of images captured by a monochrome 1.6 megapixel (1440x1080) CMOS camera (Thorlabs Zelux) with 3.45 $\mu$m square pixels and a wide-angle 8 mm f1.8 Edmund Optics lens opened to maximum aperture and focused at approximately 3 meters. This particular choice of lens and sensor leads to an optical system that is not diffraction limited despite the relatively small pixel size.
The pixels in our camera are smaller than many professional photography cameras, which are also not diffraction limited when using large apertures. Pixel shift technology that is available on various cameras could in principle be used to implement the imaging systems described here, although that would likely require special access to the control system of the camera.
For the experiments described here, instead of moving the sensor, we move the camera. The camera was mounted on a movable stage and controlled by a nanopositioning system (Nanomotion II) that uses stepper motors to move the stage at nanometer level resolution in two separately controlled axes. Movements made by the stage have high precision and repeatability, though no position readout is available. The stage is connected to and controlled by a computer via a serial port.
A static platform was mounted 3 meters away and parallel to the camera to hold targets to be imaged. The illumination source was a battery powered LED tube light (10 inch PavoTubeII 6C RGBWW).
(a) Camera and stage setup
(b) Camera is aimed at a target platform
Figure 9: (a) The camera is attached to a linear stage that adjusts the distance between the camera and the target. The linear stage is secured to the Nanomotion II stage. (b) Example of printed design and small object placed on the target platform.
We imaged a variety of targets including small three-dimensional objects (ex. Ketchup packet and LEGO(R) sets) and multiple 8.5” x 11” pages of paper containing designs printed by a standard black and white laser printer, such as a page of text and a to-scale version of the 1951 USAF resolution target.
The USAF target is a standard test chart used to determine the resolution of an optical system. It consists of vertical and horizontal sets of three bars at varying sizes and spatial frequencies that are organized in numbered groups, each with six labeled sets. Once imaged, intensity profiles across the bars can be analyzed; past the resolution limit, three distinct peaks/valleys will no longer be visible in the profile. The group and element number of the smallest resolvable set is used to look up the width and spacing of these bars, and this value, in micrometers or lp/mm, is a measure of the resolution of the system.
For each set of experiments, flat-field images were acquired by imaging a blank sheet of paper, which served as the background for all targets.
Our system is highly controlled via a stationary target, fixed lighting, and long acquisition times. One application where these conditions are available is in digital archiving. In this field, unique camera architectures have been developed to capture high-resolution images (see, e.g. [4]).
6.2 Grid Exposures
To acquire a static grid of images, the stage was moved through a series of positions in a plane parallel to the target, coming to a stop at each position and waiting for two frames to be captured, before continuing. The overall grid motion was repeated five times. Averaging many frames reduces the effects of sensor noise and running multiple trials of the grid sequence minimizes the effects of temporal illumination variation during the acquisition.
The number of acquired low-resolution images is determined by the desired resolution increase factor $f$. The actual position of these acquired images depends on the magnification factor induced by the camera optics and the distance to the target.
Moving the camera is equivalent to moving the target, and the magnification $m$ of the system relates the distance traveled by the camera to the distance traveled by the image on the sensor. We can determine $m$ using the optical properties of the camera and the distance to the target. In practice, we find $m$ by imaging an object with known length and calculating its size on the sensor by counting the number of pixels it spans and multiplying by the pixel size.
With grid capture (see Section 6.2), the image of the target moves by $\frac{\Delta}{f}$ between adjacent grid locations, for a total of $f\times f$ locations. Taking into account the magnification $m$, the distance the stage travels between neighboring grid positions, i.e. step size, is $\frac{\Delta}{fm}$.
For $f=8$, we use 64 grid locations. The magnification of our system was calculated to be 0.002484 (we use a wide-angle lens), which gives a step size of 0.1736 mm.
Due to the large field of view of the camera, our targets occupy a relatively small (150x190) area in the images acquired, which means that, for the experiments presented here, we are only using a small region of the physical sensor in the camera. In other applications, the same setup would allow for capturing very wide field of views.
Figure 10 shows several results, each a small ROI of the target area. We note that in each example we can see high-resolution details appear in the super-resolution images, such as the punctuation and to some extent serifs in the text, elements in USAF group -1 in the resolution target, and the serrated top edge of the ketchup packet.
(a) Static image (1 of 64)
(b) Wiener reconstruction
(c) TV reconstruction
Figure 10: Super-resolution using an 8x8 grid of static images. We show a small crop of various targets.
Moving the camera instead of the sensor can lead to motion parallax. Figure 11 shows that a constant magnification model can still be used for three-dimensional objects that protrude out of the target plane. The front of the LEGO(R) bus is approximately 15 cm away from the calibrated target plane and high-resolution details on the minifigures are still resolved. Even with twice as much depth variation, the parallax effect would be small due to the large distance between the camera and the object.
Note that in Figure 11, the striped overalls of a LEGO minifigure appear uniform in a static image (a) but have visible stripes in the super-resolution image (c). The stripes can be seen accurately as lines with widths 3-4x smaller than the original pixel size, indicating a 3-4x improvement.
(a) Static image (1 of 64)
(b) Wiener reconstruction
(c) TV reconstruction
Figure 11: Super-resolution using an 8x8 grid of static images. The bottom row shows a crop of the front section of the bus. Using a moving camera, instead of a moving sensor, requires the camera motion to be proportional to the distance from the object. Nonetheless, we can image three-dimensional objects that are relatively close to a target plane. In this case, a constant magnification model is still accurate.
For the examples above, we averaged multiple frames taken from each camera position to minimize the effect of measurement noise and to determine an upper limit on the reconstruction quality obtained using a physical system. Figure 12 compares a reconstruction obtained using a single frame from each position to a reconstruction obtained using ten frames from each position. While the reconstruction using a single frame from each position has more noise, it still recovers high-resolution features.
(a) 1 frame per position
(b) 10 frames per position
Figure 12: By averaging 10 frames per position, we reduce the effect of measurement noise (TV reconstructions).
6.3 Scan Motion
To acquire images with motion, the stage was moved at constant horizontal velocity over a “long” travel range corresponding to a length of many pixels. As the stage moved, five images were captured with a short random delay set before the start of each acquisition. Once the motion across the row was completed, the height of the stage was increased by sub-pixel steps (same as in grid exposures) and the motion was repeated. The total number of heights corresponds to the selected factor $f$.
Since our stage is slow, we used a relatively long exposure to capture images with motion blur. The exposure was determined by how long it takes for the image on the sensor to move by $\Delta$ (the sensor pixel size). The maximum velocity of the stage is only 2.5 mm/s and therefore, for these experiments we used an exposure time of about five hundred milliseconds.
Due to the fact that the stage has no position readout capabilities and the camera cannot be synchronized with it, the initial camera position for each acquisition must be estimated. The exposure time, vertical position, and velocity of the stage during each capture are known. The printed targets include a black rectangle, which provides a step edge that is used for the horizontal position estimation; similar to the blurred point source shown in Figure 5, the initial position of a sharp edge can be localized to sub-pixel location from a moving camera.
(a) Static image
(b) TV reconstruction
Figure 13: Super-resolution using 4 horizontal scans with constant velocity and 4 exposures per scan. (a) shows an ROI from a static image. (b) shows the super-resolution result using 16 images taken with constant velocity motion.
Figure 13 illustrates some results of the scanning motion procedure. These examples demonstrate the use of a moving camera to increase resolution beyond what is visible in a static image, despite the induced motion blur.
7 Conclusion and Future Work
We described several results related to the limits of super-resolution with a moving sensor/camera.
Our contributions include:
1) A new analysis and demonstration of super-resolution and deconvolution with a box using sparse image priors.
2) The idea and demonstration that motion blur can be beneficial for super-resolution.
(a) $|\operatorname*{FFT}(B_{4})|$ (in dB)
(b) $|\operatorname*{FFT}(B_{7})|$ (in dB)
(c) $|\operatorname*{FFT}(B_{4})|+|\operatorname*{FFT}(B_{7})|$ (in dB)
Figure 14: Bode plots of a two-dimensional box filter of width 4 (a), width 7 (b), and the sum of the two (c). The last plot illustrates how using pixels of two different sizes should eliminate most of the ambiguity for super-resolution.
We have shown that by using a grid of static images, the ambiguity in super-resolution can be restricted to relatively few Fourier coefficients. It is interesting to note that by using multiple magnifications, or a sensor with multiple pixel sizes, one could potentially remove all of the ambiguity in the super-resolution problem. Figure 14 illustrates the Fourier transform of a discrete periodic box of width 4 and similarly of a box of width 7. We see that the zeros of the two Fourier transforms occur in different places; if we were to record the convolution of a high-resolution image with boxes of two sizes, we would eliminate essentially all ambiguity in the super-resolution problem.
The idea that motion blur can be used to improve resolution is somewhat counterintuitive and leads to the notion that to take a sharper picture, one should shake or vibrate the camera. In synthetic experiments, we have shown that simply vibrating the image sensor during capture can lead to higher resolution images. In this case reconstruction involves solving a deconvolution problem while increasing resolution at the same time.